write a quadratic equation having the given numbers as solutions.
3/5, -9

Found 3 solutions by Greenfinch, JBarnum, karibet:
Answer by Greenfinch(383)

Equation is
(x + 9)(5x - 3) = 0 which expands to
5x^2 + 42x - 27 = 0

Answer by JBarnum(2044)

x=3/5, -9
x=3/5
5x=3
5x-3=0
x=-9
x+9=0
(5x-3)(x+9)
5x^2-3x+45x-27
5x^2+42x-27

Answer by karibet(1)

5(x+9)+2= -5(x-1)-1